The sum of the areas of the two squares is 325 square meters. The side of one increased by the side of the other is 25 meters. Find the length of a side each.
Answer by stanbon(75887)

The sum of the areas of the two squares is 325 square meters. The side of one increased by the side of the other is 25 meters. Find the length of a side each.
------------------
Let one of the sides be "x" ; the other side is "25-x"
EQUATION:
x^2 + (25-x)^2 = 325
x^2 + 625 -50x + x^2 = 325
2x^2 -50x + 300 = 0
x^2 -25x + 150 = 0
x = [25 +- sqrt (625-4*150)]/2
x = [25 +- 5]/2
x = [15] or x = [10]
One of the sides is 15 m the other is 10m
====================================================
Cheers,
Stan H.